OpenStudy (anonymous):

how do you calculate the standard deviation if your given the standard error

OpenStudy (anonymous):

if I remember correctly its SD = SE / n - 1 for pop SD = SE / n for sample

OpenStudy (anonymous):

so if i have this info group 1: n=9, mean= 59 and SE=7 and group 2: n=11 mean=32 and SE=10 how do i know which is populatinon or sample

OpenStudy (anonymous):

well what exactly is the question asking you to do?

OpenStudy (anonymous):

what are the two sample standard deviations

OpenStudy (anonymous):

Do it separately for each info group: standard error of mean = SD / sqroot ( N ) re arrange for SD = standard error.. * sqroot ( N )

OpenStudy (anonymous):

ignore my first formula btw

OpenStudy (anonymous):

to make sure its \[SE \times \sqrt{n}\]

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thanks